OpenStudy (anonymous):

The variable y depends on x according to the equation: x+y=x^2+y^2. Then a) dy/dx=2x+2y b) dy/dx = x+y-x^2-y^2 c) dy/dx = 1-2y/1-2x d) dy/dx = 1-2x/2y-1

OpenStudy (anonymous):

implicit differentiation \[ 1 +y^{\prime}=2x+2yy^{\prime}\] \[ 1-2x =+2yy^{\prime}-y^{\prime}\] \[ 1-2x =+(2y-1)y^{\prime}\] \[ \frac{1-2x}{(2y-1)} =y^{\prime}\]

OpenStudy (anonymous):

so D

OpenStudy (anonymous):

x+y=x^2+y^2 y=x^2+y^2 -x dy/dx=2x+2y dy/dx -1 dy/dx-2y dy/dx=2x -1 (1-2y) dy/dx=2x -1 dy/dx=(2x -1)/(1-2y) ................. D)